1/ As the assistant to the Chief Financial Oce for Volkswagen, you've been tasked with managing their

international operations so as to facilitate the coordination of their manufacturing networks. More speci-

cally, you've be presented with the following scenarios and asked to present a 'transnational solution' with

cash currently on hand. As such, you are considering exchanging Swiss Francs (CHF) for Japanese Yen

(JPY). At the bank you see the following.

CHF/USD 1.6933

USD/JPY: 115.44

What is the JPY/CHF exchange rate?

What is the CHF/JPY exchange rate?

If the CHF appreciates by 10%; what would the new JPY/CHF rate be?

If the JPY lost 25% of its quoted value against the CHF, what would the new JPY/CHF rate be?
Answer #1

What is the JPY/CHF exchange rate?

JPY/CHF exchange rate = 1 / [CHF/USD x USD/JPY] = 1 / (1.6933 x 115.44) =  0.0051

What is the CHF/JPY exchange rate?

CHF/JPY exchange rate = CHF/USD x USD/JPY = 1.6933 x 115.44 =  195.4746

If the CHF appreciates by 10%; what would the new JPY/CHF rate be?

The new JPY/CHF = (Old JPY/CHF] / (1 + appreciaton in CHF) = 0.0051 / (1 + 10%) = 0.0047

If the JPY lost 25% of its quoted value against the CHF, what would the new JPY/CHF rate be?

The new JPY/CHF = (Old JPY/CHF] x (1 - depreciaton in JPY) = 0.0051 x (1 - 25%) = 0.0038